X^3+3X^2+X+9
      /  /   /      dx
    (X^2+1)(X^2+3)
    
    
     
1
Expert's answer
2021-01-26T03:38:01-0500

"\begin{aligned}
&\int\dfrac{x^3+3x^2+x+9}{ (x^2+1)(x^2+3)} dx = \\ \\
&\textsf{performing partial fraction decomposition}\\
&\dfrac{x^3+3x^2+x+9}{ (x^2+1)(x^2+3)}= \dfrac{A}{(x²+1)} + \dfrac{B}{(x²+3)} \\ &\dfrac{x}{(x²+3)} + \dfrac{3}{(x²+1)} \\ \\ \\
&\therefore \int\dfrac{x^3+3x^2+x+9}{ (x^2+1)(x^2+3)} dx = \int\dfrac{x}{(x²+3)} dx+ \int \dfrac{3}{(x²+1)}dx




\end{aligned}"


for,

"\int \dfrac{x}{x²+3} dx \\"

let u = x² + 3, du = 2xdx

"\dfrac12\int\dfrac1udu = \dfrac{\ln u}2 + c = \dfrac12\ln (x²+3) + c"


for,

"\int \dfrac{3}{x²+1} dx = 3\int \dfrac{1}{x²+1} dx"

but, "\int \dfrac{1}{x²+1} dx = \tan^{-1}x \\"

"\int \dfrac{3}{x²+1} dx = 3\tan^{-1}x + c"




"\begin{aligned}\therefore\int\dfrac{x^3+3x^2+x+9}{ (x^2+1)(x^2+3)} dx = \dfrac12\ln (x²+3) + 3\tan^{-1}x + c
\end{aligned}"